# Thermodynamic and observational implications of black holes in toroidal geometry

### v1

@@ V Sparsity of Hawking radiations and Emission of Energy through Different Entropy Frameworks
$\eta_{\mathrm{Exp}}$ $=$ $\frac{16\mathbb{C}\pi(S-1)^{2}}{27\mathbb{g}\left\{Q^{2}-4P(S-1)\right\}^{2}}~,$
$\eta_{\mathrm{R}}$ $=$ $\frac{4\mathbb{C}\pi\alpha^{2}e^{-2\alpha S}\left(e^{\alpha S}-1\right)^{2}}{27\mathbb{g}\left\{\alpha^{2}Q^{2}-2P\left(e^{\alpha S}-1\right)^{2}\right\}^{2}}\,.$ (46)
@@
Figure 6: Sparsity $\eta$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\alpha=1$ (in Rèny entropy case) and we substitute $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve) for different curves, in Figs. (6(a))-(6(c)). However, in Fig. (6(d)), we insert Q=1,P=1Q=1,~P=1 and for red solid, blue dotted, and black dotted-dashed curves we utilize $\alpha=0.01,~0.5,~1$, respectively.
Let us mention here that in this analysis of sparsity, we employed the conjugate temperature corresponding to the entropy framework to maintain consistency. To analyze the influence of BH system parameters on Hawking sparsity using different entropy approaches, Fig. 6 presents $\eta$ as the function of entropy $S$ relationship at $P=1$ and different trajectories corresponding to charge $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $1$ (black dotted-dashed curve). We present the sparsity of Hawking radiation by using Hawking-Bekenstein entropy in Fig. 6(6(a)) and exponential corrected entropy in Fig. 6(6(b)). Similarly, we present $\eta$ as the function of entropy by putting Q=0.08,0.5,1Q=0.08,~0.5,~1 for Rènyi entropy in Fig. 6(6(c)) while in Fig. 6(6(d)) we employ different values of non-extensive parameter $\alpha=0.01,~0.5,~1$. It observes that as entropy increases, it leads to a reduction in sparsity, which is not the case for Schwarzschild BHs. Also, for smaller values of entropy, sparsity $\eta$ increases more than the Schwarzschild one, which indicates that, during this phase of evaporation, the BH emits radiation with a lower intensity than Hawking radiation. As $S$ increases, $\eta$ steadily decreases, and approaches zero in the limit. In the case of exponential corrected entropy, $\eta$ decreases as the entropy increases, but after some values of $S$, it shows constant behavior, which is quite evident in Fig. 6(6(b)). Significant changes in the decay behavior arise either due to modifications in the BH parameter space or from adopting different entropy formalisms.
@@
Figure 7: The rate of energy emission $\varepsilon_{\omega t}$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\omega=1,~\alpha=1$ (in Rèny entropy case) and for different curves we put $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve), in Figs. (7(a))-(7(c)). But for Fig. (7(c)), we put Q=1,P=1Q=1,~P=1 and $\alpha=0.01,~0.5,~1$ for red solid, blue dotted, and black dotted-dashed curves, respectively.
where $T$ is the temperature, which in our analysis corresponds to the different entropy frameworks such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies. We plot the rate of energy emission $\varepsilon_{\omega t}$ in the form of the Hawking-Bekenstein, exponential corrected, and Rènyi entropies in Fig. 7 by setting $\alpha=1$ and we obtain different trajectories for $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $Q=1$ (black dotted-dashed curve) for Figs. 7(7(a)-7(c)) while in Fig. 7(7(d)), we put $Q=1,~\omega=1,~P=1$ and $\alpha=0.01$ (red solid curve), $\alpha=0.5$ (blue dotted curve) and $\alpha=1$ ( black dotted-dashed dashed curve). For a charged torus-like BH, it is noticed that the rate of energy emission is increasing initially with a negative trend for small ranges of entropy, but as the entropy increases, the values of $\varepsilon_{\omega t}$ also increase with a positive trend. While in the case of exponential corrected entropy, we observe that the rate of emission energy is increasing for all the values of entropy with a positive trend, which we can clearly observe in Fig. 7(7(b)). However, the behavior of $\varepsilon_{\omega t}$ in Figs. 7(7(a)-7(d)) demonstrates that for large values of entropy, BH radiates permanently. The variation in the $\varepsilon_{\omega t}$ led BHs in high-energy configurations to emit radiation of low frequency with large values of entropy, which makes their identification very complicated. These phenomena are quite significant in the observational astronomy of BH evaporation for differentiating between the traditional general relativity and other gravity frameworks.
@@ VI Frequency Shift
This section transitions our focus from thermodynamic to observational analysis through a study of the frequency shift experienced by test particles orbiting charged torus-like BHs. Moreover, we derive some basic expressions that are essential in this study to examine the frequency shift, for example, we need to derive geodesics for timelike and null particles by adopting the methodology given in Ref. Martinez-Valera:2023guj in our case (charged torus-like BH) [^7] Here one can notice that the components construction in our case is significantly different from the spherically symmetric spapcetime given in Ref. Martinez-Valera:2023guj , for example in spherically symmetric spacetime we have $g_{\phi\phi}=r^{2}\sin^{2}\!\theta$ while in our case it is $g_{\phi\phi}=r^{2}$ as dicussed in Refs. Huang ; Han ; Ali:2023pyv ; Feng:2021vey ; Liang:2021elg .. First, we define the line element, which we have already provided in Eq. (3), and the metric components with respect to this line element can be written as
$g_{tt}$ $=$ $-\left(-\frac{2M}{\pi\ r}+\frac{4Q^{2}}{\pi\ r^{2}}-\frac{\Lambda\ r^{2}}{3}\right),\quad g_{rr}=\left(-\frac{2M}{\pi\ r}+\frac{4Q^{2}}{\pi\ r^{2}}-\frac{\Lambda\ r^{2}}{3}\right)^{-1},\quad g_{\theta\theta}=r^{2},$ (49)
@@ VI.1 Massive Particles
$\mathbb{E}=\frac{\mathcal{E}}{m}=-\zeta_{\mathfrak{u}}\mathbb{U}^{\mathfrak{u}}=g_{tt}\mathbb{U}^{\mathfrak{u}},\quad\mathbb{L}=\frac{\mathcal{L}}{m}=\Psi_{\mathfrak{v}}\mathbb{U}^{\mathfrak{v}}=g_{\phi\phi}\mathbb{U}^{\mathfrak{v}}\,$, (51)
@@
where the derivative with respect to radial coordinate $r$ is presented by prime . The above-mentioned condition allows us to write $\mathbb{E}$ and $\mathbb{L}$ in the form of $g_{tt}$$g_{\phi\phi}$, respectively, which takes the following shape
@@
$\mathbb{E}^{2}$ $=$ $-\frac{\left(6Mr-12Q^{2}+\pi\Lambda r^{4}\right)^{2}}{9\pi r^{2}\left(3Mr-8Q^{2}\right)}\,$, (57)
$\mathbb{L}^{2}$ $=$ $\frac{r^{2}\left(-3Mr+12Q^{2}+\pi\Lambda r^{4}\right)}{6Mr-12Q^{2}+\pi(\Lambda-3)r^{4}}\,.$ (58)
@@
$\mathbb{U}^{t}_{\mathrm{em}}=-\frac{E}{g_{tt}}=\sqrt{\frac{\pi\ r^{2}}{8Q^{2}-3Mr}}\,,\quad\mathbb{U}^{\phi}_{\mathrm{er}}=\frac{\mathbb{L}}{g_{\phi\phi}}=\sqrt{\frac{3\left(-3Mr+12Q^{2}+\pi\Lambda r^{4}\right)}{r^{2}\left(3Mr-8Q^{2}\right)}}\,.$ (59)
@@
$\frac{1}{r^{2}\left(3Mr-8Q^{2}\right)}\left\{\frac{4\left(8Q^{2}-3Mr\right)^{2}}{6Mr-12Q^{2}+\pi\Lambda r^{4}}-5Mr+16Q^{2}\right\}=0\,.$ (62)
@@ VI.2 Null Particles
By adopting a similar methodology to that adopted for spherically symmetric spacetime in Refs. Martinez-Valera:2023guj to our case (charged BH in toroidal spacetime), the energy and angular momentum of the photon particles can be written as
@@
$b_{\gamma}=\pm\sqrt{\frac{g_{\phi\phi}}{g_{tt}}}=\pm\sqrt{3\pi}\sqrt{\frac{r^{4}}{6Mr-12Q^{2}+\pi\Lambda r^{4}}}\,.$ (70)
@@
If we examine Eq. (74), the $\pm$ signs reflect the direction of mass; for instance, a positive sign indicates clockwise rotation, while a negative sign indicates counterclockwise motion. Moreover, it is irrelevant whether the rotation of the particle is clockwise or counterclockwise in the cases of spherically symmetric spacetime and axisymmetric toroidal spacetime. Now, the general expression as given in Ref. Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau ; Martinez-Valera:2023guj can be obtained by utilizing Eqs. (59), (70), and (74), which can be written as
$1+z_{\mathrm{BH_{{}_{{}_{{}_{1,2}}}}}}$ $=$ $\frac{\sqrt{g_{\phi\phi}^{\prime}}\ \pm\sqrt{\frac{g_{\phi\phi}}{g_{tt}}g_{tt}^{\prime}}}{g_{\phi\phi}g_{tt}^{\prime}-g_{\phi\phi}^{\prime}\ g_{tt}}=\frac{\sqrt{\pi}\left(\sqrt{r_{\mathrm{e}}}\pm\sqrt{\frac{3r_{\mathrm{e}}\left(8Q^{2}-3Mr_{\mathrm{e}}\right)}{6Mr_{\mathrm{e}}-12Q^{2}+\pi\Lambda r^{4}_{\mathrm{e}}}+r_{\mathrm{e}}}\ \right)}{\sqrt{\frac{8Q^{2}}{r_{\mathrm{e}}}-3M}}\,.$ (75)
@@
$z_{\mathrm{ke_{\pm}}}$ $=$ $\pm\frac{1}{\sqrt{\frac{8Q^{2}}{r_{\mathrm{e}}}-3M}}\left\{\sqrt{\frac{3r_{\mathrm{e}}\left(8Q^{2}-3Mr_{\mathrm{e}}\right)}{6Mr_{\mathrm{e}}-12Q^{2}+\pi\Lambda r^{4}_{\mathrm{e}}}+r_{\mathrm{e}}}\right\}\,.$ (77)
@@
Figure 8: The redshift $z_{1}$ (left panel) and the blueshift (right panel) in $r_{\mathrm{e}}-\Lambda$ plane by setting $M=1$ and $Q=1$. In addition, the black curve presents $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, which we numerically compute.
@@
Figure 9: Graphs of the redshift (upper left panel), blueshift (upper right panel), and gravitational shift (lower middle panel) as the function of cosmological constant $\Lambda$ by substituting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and for different trajectories we varied the electric charge $Q$ such as $Q=0.9$ (red curve), $Q=1$ (blue dotted curve) and $Q=1.3$ (black dotted-dashed curve).
@@
We graphically present the behavior of redshift (left panel) and blueshift (right panel) by putting M=1,Q=1M=1,~Q=1 in the $r_{\mathrm{e}}-\Lambda$ plane. Additionally, it’s worth noting that we select $r_{\mathrm{e}}-\Lambda$ by considering the range of $\Lambda\in[-1,1]$ and maintaining consistency in our study. Moreover, the black curve in Fig. 8 is $r_{\mathrm{ISCO}}$, which we obtain by numerically solving Eq. (62), and it also satisfies the condition of orbit stability $r_{\mathrm{ISCO}}\leq r$. We observe that in the case of a charged Torus-like BH $r_{\mathrm{ph}}$, which is obtained by using the condition provided in Ref. Martinez-Valera:2023guj , violates $r_{\mathrm{ph}}\nleq r_{\mathrm{ISCO}}$. The reason for this violation is that the metric function for charged torus-like BH is different from the Schwarzschild case for example in our case we compute $r_{\mathrm{ph}}=\frac{8Q^{2}}{3M}$ (which also agrees with the one given in Ref. Ali:2023pyv ), it indicates that $r_{\mathrm{ph}}$ is inversely in relation with the mass and directly proportional with $Q$. In the case of the Schwarzschild BHs, we have $r_{\mathrm{ph}}=3M$. Thereby, to satisfy the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$, if we take $M>1$ and $Q\leq 1$. We observe in Fig. 8 that there are opposite impacts of $r_{\mathrm{e}}$ and $\Lambda$ on redshift and blueshift frequencies, for example, as $r_{\mathrm{e}}$ and $\Lambda$ increase, the frequency of blueshift (left panel) and redshift (right panel) also decreases and increases, respectively. Consequently, we assume that these parameters can balance the frequency shift at particular points marked by grey curves, which represent a constant level of contours $z_{1,2}$.
@@
Table 1: Summary of the numerical values of $r_{\mathrm{ISCO}}$ and also its comparison with the RN-AdS and Schwarzschild BHs. One can also verify the condition violation of $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ while it is always satisfied in the case of spherically symmetric spacetime, such as RN-AdS and Schwarzschil BHs. We also mention here that CTL BH presents the abbreviation of charge torus-like BH.

### v2

@@ V Sparsity of Hawking radiations and Emission of Energy through Different Entropy Frameworks
$\eta_{\mathrm{Exp}}$ $=$ $\frac{16\mathbb{C}\pi^{4}(S-1)^{2}}{27\mathbb{g}\left(\pi^{2}Q^{2}-4P(S-1)\right)^{2}}~,$
$\eta_{\mathrm{R}}$ $=$ $\frac{4\mathbb{C}\pi^{4}\alpha^{2}e^{-2\alpha S}\left(e^{\alpha S}-1\right)^{2}}{27\mathbb{g}\left(\pi^{2}\alpha^{2}Q^{2}-2P\left(e^{\alpha S}-1\right)^{2}\right)^{2}}\,.$ (46)
@@
Figure 6: Sparsity $\eta$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\alpha=1$ (in Rèny entropy case) and we substitute $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve) for different curves, in Figs. (6(a))-(6(c)). However, in Fig. (6(d)), we insert Q=1,P=1Q=1,~P=1 and for red solid, blue dotted, and black dotted-dashed curves we utilize $\alpha=0.2,~0.6,~1$, respectively.
Let us mention here that in this analysis of sparsity, we employed the conjugate temperature corresponding to the entropy framework to maintain consistency. To analyze the influence of BH system parameters on Hawking sparsity using different entropy approaches, Fig. 6 presents $\eta$ as a function of entropy $S$ relationship at $P=1$ and different trajectories corresponding to charge $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $1$ (black dotted-dashed curve). We present the sparsity of Hawking radiation by using Hawking-Bekenstein entropy in Fig. 6(6(a)) and exponential corrected entropy in Fig. 6(6(b)). Similarly, we present $\eta$ as the function of entropy by putting Q=0.08,0.5,1Q=0.08,~0.5,~1 for Rènyi entropy in Fig. 6(6(c)) while in Fig. 6(6(d)) we employ different values of non-extensive parameter $\alpha=0.2,~0.6,~1$. It observes that as entropy increases, it leads to a reduction in sparsity, which is not the case for Schwarzschild BHs. Also, for smaller values of entropy, sparsity $\eta$ increases more than the Schwarzschild one, which indicates that, during this phase of evaporation, the BH emits radiation with a lower intensity than Hawking radiation. As $S$ increases, $\eta$ steadily decreases, and approaches zero in the limit. In the case of exponential corrected entropy, $\eta$ decreases as the entropy increases, but after some values of $S$, it shows constant behavior, which is quite evident in Fig. 6(6(b)). Significant changes in the decay behavior arise either due to modifications in the BH parameter space or from adopting different entropy formalisms.
@@
Figure 7: The rate of energy emission $\varepsilon_{\omega t}$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\omega=1,~\alpha=1$ (in Rèny entropy case) and for different curves we put $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve), in Figs. (7(a))-(7(c)). But for Fig. (7(d)), we put Q=1,P=1Q=1,~P=1 and $\alpha=0.01,~0.5,~1$ for red solid, blue dotted, and black dotted-dashed curves, respectively.
where $T$ is the temperature, which in our analysis corresponds to the different entropy frameworks such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies. We plot the rate of energy emission $\varepsilon_{\omega t}$ in the form of the Hawking-Bekenstein, exponential corrected, and Rènyi entropies in Fig. 7 by setting $\alpha=1$ and we obtain different trajectories for $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $Q=1$ (black dotted-dashed curve) for Figs. 7(7(a)-7(c)) while in Fig. 7(7(d)), we put $Q=1,~\omega=1,~P=1$ and $\alpha=0.2$ (red solid curve), $\alpha=0.6$ (blue dotted curve) and $\alpha=1$ ( black dotted-dashed dashed curve). For a charged torus-like BH, it is noticed that the rate of energy emission is increasing initially with a negative trend for small ranges of entropy, but as the entropy increases, the values of $\varepsilon_{\omega t}$ also increase with a positive trend. While in the case of exponential corrected entropy, we observe that the rate of emission energy is increasing for small values of entropy with a positive trend, and for large values of entropy, the rate of emission energy slightly decreases, and after that, it becomes constant, which we can clearly observe in Fig. 7(7(b)). However, the behavior of $\varepsilon_{\omega t}$ in Figs. 7(7(a),7(c),7(d)) demonstrates that for large values of entropy, BH radiates permanently. The variation in the $\varepsilon_{\omega t}$ led BHs in high-energy configurations to emit radiation of low frequency with large values of entropy, which makes their identification very complicated. These phenomena are quite significant in the observational astronomy of BH evaporation for differentiating between the traditional general relativity and other gravity frameworks.
Furthermore, from our analysis, it can be inferred that quantum gravitational corrections incorporated through the exponential corrected entropy suppressed strong thermal fluctuations and decreased the sparsity of Hawking radiation, effectively serving as a mechanism for information preservation via strong horizon correlation that slows down the evaporation of BHs.
@@ VI Frequency Shift
This section transitions our focus from thermodynamic to observational analysis through a study of the frequency shift experienced by test particles orbiting charged torus-like BHs. Moreover, we derive some basic expressions that are essential in this study to examine the frequency shift, for example, we need to derive geodesics for timelike and null particles by adopting the methodology given in Ref. Martinez-Valera:2023guj in our case (charged torus-like BH) [^7] Here one can notice that the components construction in our case is significantly different from the spherically symmetric spapcetime given in Ref. Martinez-Valera:2023guj , for example in spherically symmetric spacetime we have $g_{\phi\phi}=r^{2}\sin^{2}\!\theta$ while in our case it is $g_{\phi\phi}=r^{2}$ as dicussed in Refs. Huang ; Han ; Ali:2023pyv ; Feng:2021vey ; Liang:2021elg .. We also want to highlight that all physical quantities are presented in natural (planks) units, making them dimensionless and ensuring that any comparison with astrophysical data remains qualitative. First, we define the line element, which we have already provided in Eq. (3) by utilizing $\Lambda=-3/l^{2}$, and the metric components with respect to this line element can be written as
$g_{tt}$ $=$ $-\left(-\frac{2M}{\pi\ r}+\frac{4Q^{2}}{\pi\ r^{2}}+\frac{r^{2}}{l^{2}}\right),\quad g_{rr}=\left(-\frac{2M}{\pi\ r}+\frac{4Q^{2}}{\pi\ r^{2}}+\frac{r^{2}}{l^{2}}\right)^{-1},\quad g_{\theta\theta}=r^{2},$ (49)
@@ VI.1 Massive Particles
$\mathbb{E}=\frac{\mathcal{E}}{m}=-\zeta_{\mathfrak{u}}\mathbb{U}^{\mathfrak{u}}=-g_{tt}\mathbb{U}^{\mathfrak{u}},\quad\mathbb{L}=\frac{\mathcal{L}}{m}=\Psi_{\mathfrak{v}}\mathbb{U}^{\mathfrak{v}}=g_{\phi\phi}\mathbb{U}^{\mathfrak{v}}\,$, (51)
@@
where prime shows the derivative with respect to the radial coordinate $r$. The above-mentioned condition allows us to write $\mathbb{E}$ and $\mathbb{L}$ in the form of $g_{tt}$$g_{\phi\phi}$, respectively, which takes the following shape
@@
$\mathbb{E}^{2}$ $=$ $-\frac{\left(l^{2}\left(4Q^{2}-2Mr\right)+\pi r^{4}\right)^{2}}{\pi l^{4}r^{2}\left(3Mr-8Q^{2}\right)}\,$, (57)
$\mathbb{L}^{2}$ $=$ $\frac{l^{2}r^{2}\left(Mr-4Q^{2}\right)+\pi r^{6}}{l^{2}\left(-2Mr+4Q^{2}+\pi r^{4}\right)+\pi r^{4}}\,.$ (58)
@@
$\mathbb{U}^{t}_{\mathrm{em}}=-\frac{\mathbb{E}}{g_{tt}}=\sqrt{\frac{r^{2}\pi}{8Q^{2}-3Mr}}\,,\quad\mathbb{U}^{\phi}_{\mathrm{er}}=\frac{\mathbb{L}}{g_{\phi\phi}}=\sqrt{\frac{l^{2}\left(Mr-4Q^{2}\right)+\pi r^{4}}{l^{2}r^{2}\left(8Q^{2}-3Mr\right)}}\,.$ (59)
@@
$-\frac{2\left(l^{2}\left(6M^{2}r^{2}-36MQ^{2}r+64Q^{4}\right)+3\pi r^{4}\left(5Mr-16Q^{2}\right)\right)}{r^{2}\left(3Mr-8Q^{2}\right)\left(l^{2}\left(4Q^{2}-2Mr\right)+\pi r^{4}\right)}=0\,.$ (62)
@@ VI.2 Null Particles
By adopting a similar methodology to that adopted for spherically symmetric spacetime in Ref. Martinez-Valera:2023guj to our case (charged BH in toroidal spacetime), the energy and angular momentum of the photon particles can be written as
@@
$b_{\gamma}=\pm\sqrt{\frac{g_{\phi\phi}}{g_{tt}}}=\pm\sqrt{\pi}\sqrt{\frac{l^{2}r^{4}}{l^{2}\left(4Q^{2}-2Mr\right)+\pi r^{4}}}\,.$ (70)
@@
If we examine Eq. (74), the $\pm$ signs reflect the direction of mass; for instance, a positive sign indicates clockwise rotation, while a negative sign indicates counterclockwise motion. Moreover, it is irrelevant whether the rotation of the particle is clockwise or counterclockwise in the cases of spherically symmetric spacetime and axisymmetric toroidal spacetime. Now, the general expression as given in Refs. Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau ; Martinez-Valera:2023guj can be obtained by utilizing Eqs. (59), (68), and (74), which can be written as
$1+z_{\mathrm{BH_{{}_{{}_{{}_{1,2}}}}}}$ $=$ $\frac{\sqrt{g_{\phi\phi}^{\prime}}\ \pm\sqrt{\frac{g_{\phi\phi}}{g_{tt}}g_{tt}^{\prime}}}{g_{\phi\phi}g_{tt}^{\prime}-g_{\phi\phi}^{\prime}\ g_{tt}}=\sqrt{\pi}\Bigg(\sqrt{\frac{r^{2}_{\mathrm{e}}}{8Q^{2}-3Mr_{\mathrm{e}}}}\pm\sqrt{\frac{l^{2}\left(Mr_{\mathrm{e}}-4Q^{2}\right)+\pi r^{4}_{\mathrm{e}}}{l^{2}r^{2}_{\mathrm{e}}\left(8Q^{2}-3Mr_{\mathrm{e}}\right)}}$ (75)
$\times$ $\sqrt{\frac{l^{2}r^{4}_{\mathrm{e}}}{l^{2}\left(4Q^{2}-2Mr_{\mathrm{e}}\right)+\pi r^{4}_{\mathrm{e}}}}\Bigg)\,.$
@@
$z_{\mathrm{ke_{\pm}}}$ $=$ $\pm\sqrt{\pi}\sqrt{\frac{l^{2}r^{4}_{\mathrm{e}}}{l^{2}\left(4Q^{2}-2Mr_{\mathrm{e}}\right)+\pi r^{4}_{\mathrm{e}}}}\sqrt{\frac{l^{2}\left(Mr_{\mathrm{e}}-4Q^{2}\right)+\pi r^{4}_{\mathrm{e}}}{l^{2}r^{2}_{\mathrm{e}}\left(8Q^{2}-3Mr_{\mathrm{e}}\right)}}\,.$ (77)
@@
Figure 8: The redshift $z_{1}$ (left panel) and the blueshift (right panel) in $r_{\mathrm{e}}-l$ plane by setting $M=1$ and $Q=1$. In addition, the black curve presents $r_{\mathrm{e}}=r_{\mathrm{ISCO}}$, which we numerically compute.
@@
Figure 9: Graphs of the redshift (upper left panel) and blueshift (upper right panel) as the function of $l$ by substituting $M=1,~r_{\mathrm{e}}=r_{\mathrm{ISCO}}$ and for different trajectories we varied the electric charge $Q$ such as $Q=0.8$ (red curve), $Q=0.9$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve). We also mention that since the gravitational shift (lower panel) is free from the AdS radii $l$, we demonstrate $z_{\mathrm{g}}$ in terms of $r_{\mathrm{e}}$.
@@
We graphically present the behavior of redshift (left panel) and blueshift (right panel) by putting M=1,Q=1M=1,~Q=1 in the $r_{\mathrm{e}}-l$ plane. Additionally, it’s worth noting that we select $r_{\mathrm{e}}-l$ to cover the range $l\in[0,1]$ and maintain consistency with our study. Moreover, the black curve in Fig. 8 is $r_{\mathrm{ISCO}}$, which we obtain by numerically solving Eq. (62), and it also satisfies the condition of orbit stability $r_{\mathrm{ISCO}}\leq r$. We observe that in the case of a charged torus-like BH $r_{\mathrm{ph}}$, which is obtained by using the condition provided in Ref. Martinez-Valera:2023guj , violates $r_{\mathrm{ph}}\nleq r_{\mathrm{ISCO}}$. The reason for this violation is that the metric function for charged torus-like BH is different from the Schwarzschild case for example in our case we compute $r_{\mathrm{ph}}=\frac{8Q^{2}}{3M}$ (which also agrees with the one given in Ref. Ali:2023pyv ), it indicates that $r_{\mathrm{ph}}$ is inversely in relation with the mass and directly proportional with $Q$. In the case of the Schwarzschild BHs, we have $r_{\mathrm{ph}}=3M$. Thereby, to satisfy the condition $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$, if we take $M\geq 1$ and $Q\leq 1$. We observe in Fig. 8 that there are opposite impacts of $r_{\mathrm{e}}$ and $l$ on redshift and blueshift frequencies, for example, as $r_{\mathrm{e}}$ and $l$ increase, the frequency of blueshift (left panel) and redshift (right panel) also increases (with negative behavior) and decreases (with positive behavior), respectively. Consequently, we assume that these parameters can balance the frequency shift at the points marked by the grey curves, which represent constant contour levels $z_{1,2}$.
@@
Table 1: Summary of the numerical values of $r_{\mathrm{ISCO}}$ and also its comparison with the RN-AdS and Schwarzschild BHs. One can also verify the condition violation of $r_{\mathrm{ph}}<r_{\mathrm{ISCO}}$ for $Q>1$, while we compared it in the case of spherically symmetric spacetime, such as RN-AdS and Schwarzschild BHs. We can clearly observe that the violation which we highlighted is also observed in some cases of RN-AdS BHs. We also mention here that CTL BHs present the abbreviation of charge torus-like BH.